OpenStudy (lukecrayonz):

The sum of x and 4 is the product of 10 and 8.

OpenStudy (lukecrayonz):

4+x=10*8?

OpenStudy (shubhamsrg):

yep..the very same i'd guess..

OpenStudy (lukecrayonz):

@im2bad

OpenStudy (lukecrayonz):

A question I just got to: The numbers in the set x|x=a/b where a, b are integers and b=/=0, are called.. haha, you posted it like a minute ago, nice job on looking into the future.

OpenStudy (anonymous):

its the set of rational numbers. recall that rational numbers are numbers which can be represented as a ratio of two integers

OpenStudy (anonymous):

Use the Force, Luke!